# Influence of finite-temperature effects on CMB power spectrum

### v1

@@ 3.2 Analysis by statistical and machine learning techniques
The current situation, despite appearing somewhat unnatural with $\Omega_{\Lambda_{2}}\sim 10^{-8}$, may not necessarily represent an insurmountable obstacle. Instead, it could reflect a situation where the issue might be resolved by adopting a different renormalization scheme. The key feature of this alternative scheme is that the masses are adjusted in such a way that the mass-dependent terms in the potential (as shown in eq. (21)), which were neglected in the previous renormalization scheme applied to the CC problem, are kept small but not negligibly so. This should lead to a slower convergence in the large-temperature expansion since the order of $\frac{m}{T}$, where $m$ denotes the value of the renormalized mass of, say, a Higgs particle, is larger than before. This slow convergence, which was also expected in the proposed CC resolution in [5, 6], could, in turn, suggest that the renormalized cosmological constant $\Lambda_{ren}$ should be set to a value smaller than the observed cosmological constant $\Lambda_{obs}$, by several orders of magnitude. This contrasts with the approach used in the CC problem resolution, where the renormalized and observed values were taken to be on the same order.
@@ 4 Conclusion
In a flat background, it is a textbook result that finite-temperature effects generate vacuum energy terms. It is natural to expect that this result will be generalized to a curved spacetime, leading to cosmological “constant” terms. This expectation has been verified [7] within the framework of the foliation-based quantization of gravity [8, 9, 11]. The quantum corrections modify, among other things, the cosmological constant term. This work demonstrates the significant impact of the shifted cosmological constant and their implications for the broader set of cosmological parameters. Our analysis suggests that perturbative quantum gravitational effects must be accounted for in periods around recombination, as these effects are not negligible. More specifically, through the introduction of additional parameters, such as $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$, we have extended the standard $\Lambda$CDM model, offering a more sophisticated approach to fitting cosmological data. By applying the CLASS and utilizing both brute-force scans and machine learning regression techniques, we have demonstrated how the temperature-corrected cosmological parameters can lead to a more accurate fit to the Planck 2018 data. The examination of the 7- and 8- parameter models based on quartic regression has revealed that the presence of a small nonzero value of $\Omega_{\Lambda_{2}}$ has a significant impact and has shown promise in refining the predictions of the model while preserving its generalizability.
@@
While the results presented here reinforce the utility of finite-T QFT effects in cosmological studies, they also open up avenues for future research. One potential improvement can be seen from the residual plots in Figs. 5 and 6 that display certain patterns for both models, particularly for the 8-variable model. While the residuals remain centered around zero, the spread indicates minor deviations from ideal predictions, especially for higher distance values. These residual patterns suggest that while the models are highly accurate, there is room for improvement in capturing some higher-order non-linearities or interactions present in the data. Various different machine learning techniques, including artificial neural network, have been and are currently being tried. So far, the quartic regression and artificial neural network methods turn out to be best-performing. The result of the artificial neural network [23] shares the same qualitative features as those obtained here. Some complementary aspects, such as feature importance, will be reported therein.
@@
The higher order terms may well play important roles. This anticipation is based on the slow convergence anticipation as was the case in [5, 6] in which the CC problem was tackled. Refining the approach to incorporate higher-order corrections and extending the machine learning methods for enhanced predictive accuracy will thus be important next steps. In particular, it will be worth including one additional parameter, the one associated with the linear temperature term $\sim\frac{1}{a}$, and examine its and its effects on the rest of the parameters. [^8] The ellipsis in eq. (23) include the higher-order loop corrections as well as the high-orders terms in $\frac{m}{T}$ expansion. Analyzing the former terms would require even greater effort. The exponential growth in the complexity of the parameter space poses substantial challenges that demand supercomputers with greater speed and efficiency. We will report progress on some of these issues elsewhere.

### v2

@@ 3.2.2 The results with bootstrap resampling
The current situation, despite appearing somewhat unnatural with $\Omega_{\Lambda_{2}}\sim 10^{-8}$, may not necessarily represent an insurmountable obstacle. Instead, it could reflect a situation where the issue might be resolved by adopting a different renormalization scheme. The key feature of this alternative scheme is that the masses are adjusted in such a way that the mass-dependent terms in the potential of eq. (22), which were neglected in the previous renormalization scheme applied to the CC problem, are kept small but not negligibly so. This should lead to a slower convergence in the large-temperature expansion since the order of $\frac{m}{T}$, where $m$ denotes the value of the renormalized mass of, say, a Higgs particle, is larger than before. This slow convergence, which was also expected in the proposed CC resolution in [5, 6], could, in turn, suggest that the renormalized cosmological constant $\Lambda_{ren}$ should be set to a value smaller than the observed cosmological constant $\Lambda_{obs}$, by several orders of magnitude. This contrasts with the approach used in the CC problem resolution, where the renormalized and observed values were taken to be on the same order.
@@ 4 Conclusion
[…]
While the results presented here reinforce the utility of finite-$T$ QFT effects in cosmological studies, they also open up avenues for future research. A natural next step would be to repeat the full Bayesian parameter estimation using likelihood functions from the Planck 2018 analysis. This more standard approach to parameter inference - though computationally intensive - would provide an important cross-check of our results and help quantify uncertainties more systematically. As noted in the main body, this analysis was not pursued in the present work due to its high computational cost, which would have required access to more powerful computing resources.
@@
One potential improvement can be seen from the residual plots in Figs. 4 and 5, which display certain patterns for both models, particularly for the 8-variable model. While the residuals remain centered around zero, the spread indicates minor deviations from ideal predictions, especially for higher distance values. These residual patterns suggest that while the models are highly accurate, there is room for improvement in capturing some higher-order non-linearities or interactions present in the data. Various different machine learning techniques, including artificial neural networks, have been and are currently being tried. So far, the quartic regression and artificial neural network methods turn out to be best-performing. The result of the artificial neural network [27] shares the same qualitative features as those obtained here. Some complementary aspects, such as feature importance, are reported therein.
@@
The higher-order terms may well play important roles - an expectation motivated by the slow convergence observed in previous studies of the cosmological constant problem [5, 6]. Refining the current approach to incorporate these higher-order corrections, as well as extending machine learning methods for enhanced predictive accuracy, will thus be important next steps. In particular, it will be worth including an additional parameter associated with the linear temperature term, $\sim\frac{1}{a}$, and examining its impact on the rest of the parameters. [^12] The ellipsis in eq. (24) includes both higher-order loop corrections and high-order terms in the $\frac{m}{T}$ expansion. Analyzing the former terms would require significantly greater effort. While our present numerical study indicates that leading-order finite-temperature effects alone do not resolve the tension, it remains to be seen whether this further extended model could potentially ameliorate it. In contrast, see, e.g., [28] for a study in which a novel renormalization scheme achieves a similar goal. Given the exponential growth in the complexity of the parameter space, addressing these higher-order effects and thoroughly exploring their implications - especially with more complete Bayesian methods - will likely require supercomputing resources. We will report progress on these issues elsewhere.
Appendix A Onshell thermodynamics vs. offshell 1PI analysis
In section 2.2, a potential redundancy of $\Omega_{\Lambda_{2}}$ (and $\Omega_{\Lambda_{3}}$) was addressed. Nonetheless, a lingering question may ramain, given that thermodynamic relations, such as $p=\frac{1}{3}\rho$, can also be derived in a quantum field-theoretic thermodynamic framework by introducing the Helmholtz free energy. Such a framework, while useful in certain contexts, is not entirely suitable for the present discussion, as we will explain.
First, let us briefly review the framework. Consider again the Einstein-scalar system in (21):
$S_{\text{ren}}=-\int d^{4}x\left[\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi+\frac{1}{2}m^{2}\phi^{2}+\frac{\lambda}{4!}\phi^{4}\right].$ (A.1)
We define
$V_{\text{ren}}\equiv\frac{1}{2}m^{2}\phi^{2}+\frac{\lambda}{4!}\phi^{4},\quad m_{\phi}^{2}\equiv m^{2}+\frac{1}{2}\lambda\phi^{2}.$ (A.2)
The one-loop effective potential at finite temperature can be written as (see, e.g., [29] and the review in [4])
$V_{\text{one-loop}}(\phi,T)=V_{\text{ren}}+\hbar J(m_{\phi},T),\quad\text{where}\quad J(m_{\phi},T)=\int\sum_{n}\frac{1}{2}\ln(K^{2}+m_{\phi}^{2}).$ (A.3)
This leads to the expression
$J(m_{\phi},T)=-\frac{\pi^{2}}{90}T^{4}\!+\!\frac{m_{\phi}^{2}}{24}T^{2}\!-\!\frac{1}{12\pi}m_{\phi}^{3}T\!-\!\frac{m_{\phi}^{4}}{2(4\pi)^{2}}\left[\ln\left(\frac{\bar{\mu}e^{\gamma}}{4\pi T}\right)\!+\!\frac{1}{2\epsilon}\right]\!+\!\frac{\zeta(3)}{3(4\pi)^{4}}\frac{m_{\phi}^{6}}{T^{2}}+\dots$ (A.4)
where $\ln\mu^{2}=\ln(4\pi\mu^{2})-\gamma$ is the renormalization scale, $\gamma$ is the Euler constant, $\epsilon$ is the regularization parameter, and $\zeta(3)$ is the Riemann zeta function. Note that $\hbar$ is explicitly displayed. In the framework under consideration, $V_{\text{one-loop}}(\phi,T)$ is identified as the Helmholtz free energy density, and thermodynamic relations such as $p=\frac{1}{3}\rho$ can be derived for a massless case by setting the free energy density $f=\frac{\pi^{2}}{90}T^{4}$.
However, there are several important issues with applying this derivation directly to the cosmological context:
  1. 1.
    Quantum vs. Classical Thermodynamics: As seen in (A.3), $J(m_{\phi},T)$ represents a quantum correction, in contrast to the classical kinetic theory derivation of thermodynamic relations such as $p=\frac{1}{3}\rho$. The kinetic theory treatment is purely classical and does not involve quantum corrections.
  2. 2.
    On-shell Condition: For the derivation of $p=\frac{1}{3}\rho$, the scalar field $\phi$ must be set to zero, i.e., $\phi=0$, which is a trivial solution of the field equation. This implies an on-shell analysis, where the field $\phi$ is fixed to a solution. However, this is not the objective of the present work. The current analysis keeps the $\phi$-dependent part in order to obtain the quantum-corrected offshell aciton, namely the 1PI action, and focuses on the $\phi$-independent contributions to the cosmological constant. Thus, the field-theoretic derivation of $p=\frac{1}{3}\rho$ in the quantum field-theoretic thermodynamics does not imply that, for example, the $T^{4}$ contribution from vacuum diagrams is redundant to the radiation contribution.
  3. 3.
    Massive Particles in the Standard Model: In the Standard Model (SM), there are massive particles that cannot be regarded as radiation. These particles contribute to the $T^{4}$- and $T^{2}$-terms in the cosmological constant, even though they do not fit within the classical radiation picture (i.e., $p=\frac{1}{3}\rho$).
[…]
On the other hand, for matter components like dust (e.g., stars, galaxies), a purely classical treatment is appropriate. These objects are not described by Feynman diagrams but rather by kinetic theory. Hence, the only sensible approach to describing them is via kinetic theory or hydrodynamics. Since it is desirable to treat all matter components uniformly, a classical kinetic theory treatment of all matter components is the appropriate setup.
To further clarify the distinction, consider a slightly different, yet conceptually simpler approach. This approach, while arriving at the same starting point as before within the typical approximations used in cosmology, begins by formulating an effective field theory for gravity and SM particles, without including the kinetic theory sector. (The term ”effective” refers here to an effective field theory derived by keeping only the leading terms in the derivative expansion of the 1PI action.) This effective action accounts for quantum corrections, including a $T^{4}$ term (and a $T^{2}$ term) in the CC sector. Subsequently, a kinetic or hydrodynamic system - which includes matter compoenents such as $\rho_{r}$ or $\rho_{cdm}$ - can be coupled to this effective field theory. For a purely hydrodynamic system, all SM fields can be set to zero.
Appendix B Statistical glossary
Regression:
A statistical technique used to model and analyze the relationship between a dependent variable and one or more independent variables (also called features or predictors). The goal is to estimate how changes in the independent variables influence the dependent variable. Common types include linear regression, where the relationship is modeled as a straight line, and polynomial regression, where higher-degree terms allow modeling of non-linear patterns.
$R^{2}$ (Coefficient of determination):
A statistical measure that represents the proportion of variance in the dependent variable that is predictable from the independent variables. It ranges from 0 to 1, with values closer to 1 indicating better model performance.
MSE (Mean squared error):
A measure of the average squared difference between the observed actual outcomes and the outcomes predicted by the model. It is given by
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2},$
where $y_{i}$ is the actual value, $\hat{y}_{i}$ is the predicted value, and $n$ is the number of observations.
Standardization:
A preprocessing technique used to scale features so they have a mean of 0 and a standard deviation of 1. This ensures that features contribute equally to the model, especially important for distance-based and regularized methods.
Residuals:
The difference between observed values and the values predicted by a regression model. For a given observation, the residual is $e_{i}=y_{i}-\hat{y}_{i}$. Residuals are used to diagnose the fit of a regression model.
Learning curve:
A plot showing how a model’s performance (typically accuracy or error) on the training and validation datasets changes as the training set size increases. It helps assess whether adding more data improves model performance.
Overfitting:
A modeling error that occurs when a regression model is too complex and captures noise rather than the underlying pattern in the data. It usually results in high training accuracy but poor generalization to new data.
Cross-validation:
A statistical method used to estimate the performance of a model on an independent dataset. In $k$-fold cross-validation, the dataset is divided into $k$ subsets; the model is trained on k1k-1 of them and tested on the remaining subset. This process is repeated $k$ times, each time with a different test set.